OpenStudy (anonymous):

The sum of the quotients of 25x25x and the quotient of 8x2x is

OpenStudy (wintersuntime):

5x+8x=x(5+8)=13x

OpenStudy (wintersuntime):

I'm glad to help

OpenStudy (anonymous):

Oh thank you so much!